# Faithful global convergence for the rescaled Consensus–Based Optimization

### v1

@@ 2.1. Moments bounds
Let $\rho_{*}^{\alpha}\in\mathscr{P}_{2}({\mathbb{R}}^{d})$ be the unique invariant probability measure to the rescaled CBO (1.4). Then it holds that
@@
The existence has been proven in [huang2024self, Proposition 3.4]. Now we show that (2.7) holds. According to [zhang2023existence, Lemma 2.7], for any $|\varphi(x)|\leq C(1+|x|)$ it holds that
@@
Indeed, for the latter result to hold, one needs [zhang2023existence, assumption (H1) with $r_{1}=1$] whose validity in our context has already been verified in [huang2024self, Appendix A.2].
@@ 2.2. Harnack’s inequality
$A:=(a^{ij})_{i,j},\quad\text{ such that }\quad a^{ij}(x,\mu):=\frac{\sigma^{2}}{2}(\delta+|(x-\kappa\,\mathfrak{m}_{\alpha}(\mu))_{i}|)^{2}\delta_{ij}$
and $b:=(b^{i})_{i},\quad\text{ such that }\quad b^{i}(x,\mu):=-\lambda(x-\kappa\,\mathfrak{m}_{\alpha}(\mu))_{i}$
$\delta_{ij}$ being the Kronecker symbol (not to be confused with the positive constant $\delta$). In particular, when evaluated in the invariant probability measure $\rho_{*}^{\alpha}$, one obtains smooth coefficients (in $\mathcal{C}^{\infty}({\mathbb{R}}^{d})$)
@@
Moreover, the matrix $A_{*}$ is positive–definite. This identification is possible since the dependence of $A$ and $b$ on the measure is non-local – that is, it involves a quantity of the form $\int_{{\mathbb{R}}^{d}}K(y)\mbox{\rm d}\mu(y)$ which is moreover constant in $x$ (since the kernel $K$ does not involve $x$)–, rather than a pointwise (local) evaluation of (the density of) the measure $\mu$. In other words, regularity of (the density of) the measure $\mu$ does not influence regularity of the coefficients $A$ and $b$.
@@
The following result is well-known and can be found for example in [bogachev2022fokker].
@@
First observe that the coefficients $A_{\nu}$ and $b_{\nu}$ are smooth, and $A_{\nu}$ is positive–definite. Then, existence and regularity of the density is [bogachev2022fokker, Corollary 1.6.7]. Harnack’s inequality is [bogachev2022fokker, Corollary 1.7.2]. ∎
@@
Existence and uniqueness of an invariant probability measure $\rho_{*}^{\alpha}$ is obtained in [huang2024self, Proposition 3.4 & Corollary 3.8], that is, $\rho_{*}^{\alpha}$ is the unique solution to the (non-linear) equation $L_{A,b,\mu}^{*}\mu=0$ in the class of probability measures. Moreover, (2.7) ensures that $\rho_{*}^{\alpha}$ has a finite second moment. Therefore, we can consider the diffusion (linear) operator $L_{A,b,*}$ and we are in the setting of Theorem 2.8 which gives the desired result when we choose the solution to $L_{A,b,*}^{*}\mu=0$ within the probability measures. Indeed in this case, the unique probability measure $\mu$ satisfying $L_{A,b,*}^{*}\mu=0$ is $\rho_{*}^{\alpha}$. Finally, the sets $U$ and $K$ in Theorem 2.8 are chosen as the closed balls $B_{2r}(x_{0})$ and $B_{r}(x_{0})$ respectively. ∎
@@
3. Global convergence
We are now ready to state and prove our main results. We shall first treat the case where the function $f(\cdot)$ to be minimized has a unique minimizer. Then we expound the more general situation where $f(\cdot)$ has multiple minimizing regions.

### v2

@@ 2. Useful estimates
Suppose $f(\cdot)$ satisfies Assumption 2.1. Then for all $R>0$, there exists some constant $L_{\mathfrak{m}}>0$ depending on $R$ such that
(2.9) $|\mathfrak{m}_{\alpha}(\mu)-\mathfrak{m}_{\alpha}(\nu)|\leq L_{\mathfrak{m}}\mathbb{W}_{2}(\mu,\nu)\quad\forall(\mu,\nu)\in\mathscr{P}_{2,R}({\mathbb{R}}^{d})\times\mathscr{P}_{2}({\mathbb{R}}^{d})\,.$
3. On the invariant measure
This section is devoted to the limit $t\to\infty$. We will first prove existence of an invariant measure for (1.6) in §3.1. Then we will we will establish in §3.2 a $\mathbb{W}_{2}$–exponential contraction for the convergence of the law of (1.6) towards the invariant measure. This shall be next used in §3.3 where we will provide a uniform-in-$\alpha$ second moment of the invariant measure (that will be later needed in Section 4). Finally, we will show that the invariant measure verifies a Harnack type inequality which will lead to a positive and uniform-in-$\alpha$ lower bound on its support: a key result for Laplace’s principle in Section 4.
3.1. Existence
The following result ensures existence of an invariant measure for the process (1.6).
Theorem 3.1.
Suppose $f(\cdot)$ satisfies Assumption 2.1. For sufficiently small $\kappa\in(0,1)$, and for a large $\lambda>0$, there exists an invariant measure $\rho^{\alpha}_{*}\in\mathscr{P}_{2,R}({\mathbb{R}}^{d})$, for some sufficiently large $R$ that depends on $\kappa,\delta,\lambda,\sigma,b_{1},b_{2}$ and the initial distribution.
Proof.
The proof relies on [46, Theorem 2.2]. See Appendix A. ∎
Remark 3.2.
As can be seen in Appendix A, the precise condition on the parameters is explicitly given by (A.3) that is
$2\lambda-4>\lambda\,\kappa\,(1+b_{2})+4\,\kappa^{2}b_{2}$
where we recall $b_{2}$ is defined in (2.2). A simple example of sufficient conditions is when
$\kappa<\frac{1}{2(1+b_{2})}\qquad\text{ and }\qquad\lambda>4.$
3.2. Long-time limit
The next result provides a $\mathbb{W}_{2}$–exponential contraction for the evolution of the distribution of $\overline{X}$ as defined in (1.6) and cements its long-time limit.
Theorem 3.3.
Let us consider (1.6) whose distribution is $\rho^{\alpha}_{t}:=\operatorname{Law}(\overline{X}_{t})$ for all $t\geq 0$. Suppose $f(\cdot)$ satisfies Assumption 2.1. Then for sufficiently small $\kappa\in(0,1)$ and $\sigma>0$, there exists $\Theta>0$ such that
(3.1) $\mathbb{W}_{2}(\rho^{\alpha}_{t},\rho^{\alpha}_{*})^{2}\leq e^{-\Theta\,t}\,\mathbb{W}_{2}(\rho^{\alpha}_{0},\rho^{\alpha}_{*})^{2}\quad\forall\,t\geq 0$
and $\Theta$ depends on $\lambda,\sigma,\kappa$ and $L_{\mathfrak{m}}$. In particular, the invariant measure $\rho^{\alpha}_{*}$ of (1.6) is unique.
We refer to Remark 3.6 below for a precise definition of the smallness of the parameters.
Remark 3.4.
The standard CBO (1.5) (i.e. when $\kappa=1$ and $\delta=0$ in (1.6)) admits infinitely many invariant measures: see [33, Remark 1.1].
An immediate consequence of Theorem 3.3 and Lemma 2.8 is the following corollary.
Corollary 3.5.
In the situation of Theorem 3.3, it holds
$|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|\leq L_{\mathfrak{m}}e^{-\frac{\Theta}{2}\,t}\,\mathbb{W}_{2}(\rho^{\alpha}_{0},\rho^{\alpha}_{*})\quad\forall\,t\geq 0.$
In particular, one has $\lim\limits_{t\to\infty}\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})=\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})$.
Proof of Theorem 3.3.
Let us consider two processes governed by the same dynamics, but starting from different initially distributed positions. The first one is (1.6) which we recall is
$\text{with }\;\rho^{\alpha}_{t}\,=\,\text{Law}({\overline{X}}_{t})\quad\text{ and }\quad{\overline{X}}_{0}\sim\rho^{\alpha}_{0}.$
We also have, thanks to Theorem 2.6, $\operatorname{Law}(X_{t})=\rho^{\alpha}_{t}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ for all $t\geq 0$.
The second process is the stationary counter-part of (3.2), whose initial condition has the invariant measure as distribution, and that is
(3.3) ${\rm d}\overline{Z}_{t}=$ $-\lambda\left(\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\right)\,{\rm d}t\,+\,\sigma\left(\delta\,\mathds{I}_{d}+D\left(\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\right)\right)\,{\rm d}B_{t}$
$\text{with }\;\rho^{\alpha}_{*}\,=\,\text{Law}(\overline{Z}_{t})\quad\forall\,t\geq 0.$
Indeed, by definition of $\rho^{\alpha}_{*}$ being an invariant measure for (3.2), the law of $\overline{Z}_{t}$ remains equal to $\rho^{\alpha}_{*}$ for all $t\geq 0$, hence the process is stationary.
We have
${\rm d}(\overline{X}_{t}-\overline{Z}_{t})$ $=-\lambda\bigg((\overline{X}_{t}-\overline{Z}_{t})-\kappa\big(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\big)\bigg)\,{\rm d}t$
$\quad\quad\quad+\sigma\bigg(D(\overline{X}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}))-D(\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}))\bigg)\,{\rm d}B_{t}.$
As in the proof of Theorem 2.6, we use the function $h:\mathbb{R}^{d}\to\mathbb{R}$, such that $h(z)=\frac{1}{2}|z|^{2}$, and we apply Itô’s formula (2.5) which leads to
${\rm d}h(\overline{X}_{t}-\overline{Z}_{t})=$ $\bigg\{-\lambda(\overline{X}_{t}-\overline{Z}_{t})\cdot\big[(\overline{X}_{t}-\overline{Z}_{t})-\kappa(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}))\big]$
$\quad+\frac{\sigma^{2}}{2}\,\text{trace}\bigg[\bigg(D(\overline{X}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}))-D(\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}))\bigg)^{2}\bigg]\bigg\}\,{\rm d}t$
$\quad\quad+\sigma(\overline{X}_{t}-\overline{Z}_{t})\cdot\bigg[D(\overline{X}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}))-D(\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}))\bigg]\,{\rm d}B_{t}$
$=:\mathbf{F}_{t}\,{\rm d}t+\mathbf{G}_{t}\,{\rm d}B_{t}.$
The same arguments as in the beginning of the proof of [33, Theorem 2.4] (see in particular the proof of [33, equation (2.6)]), allow us to write
$\frac{{\rm d}}{{\rm d}t}{\mathbb{E}}\left[|\overline{X}_{t}-\overline{Z}_{t}|^{2}\right]=2\,{\mathbb{E}}\left[\mathbf{F}_{t}\right]$
$\quad=-2\,\lambda\,{\mathbb{E}}\left[|\overline{X}_{t}-\overline{Z}_{t}|^{2}\right]+2\,\lambda\,\kappa\,{\mathbb{E}}\left[(\overline{X}_{t}-\overline{Z}_{t})\cdot\big(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\big)\right]$
$\quad\quad\quad\quad+\sigma^{2}\,{\mathbb{E}}\left[\text{trace}\bigg(\big(D(\overline{X}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}))-D(\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}))\big)^{2}\bigg)\right].$
We can now estimate each of the terms above. We have
${\mathbb{E}}\left[(\overline{X}_{t}-\overline{Z}_{t})\cdot\big(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\big)\right]\leq\frac{1}{2}{\mathbb{E}}\left[|\overline{X}_{t}-\overline{Z}_{t}|^{2}\right]+\frac{1}{2}|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|^{2}$
and
${\mathbb{E}}\left[\text{trace}\bigg(\big(D(\overline{X}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}))-D(\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}))\big)^{2}\bigg)\right]$
$\quad=\sum\limits_{k=1}^{d}{\mathbb{E}}\left[\bigg(|\{\overline{X}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})\}_{k}|-|\{\overline{Z}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\}_{k}|\bigg)^{2}\right]$
$\quad\leq\sum\limits_{k=1}^{d}{\mathbb{E}}\left[\left|\big\{\overline{X}_{t}-\overline{Z}_{t}\big\}_{k}-\kappa\,\{\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\big\}_{k}\right|^{2}\right]$
$\quad\leq 2\sum\limits_{k=1}^{d}{\mathbb{E}}\left[\left|\big\{\overline{X}_{t}-\overline{Z}_{t}\big\}_{k}\right|^{2}+\kappa^{2}\,\left|\{\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})\big\}_{k}\right|^{2}\right]$
$\quad\leq 2\,{\mathbb{E}}\left[|\overline{X}_{t}-\overline{Z}_{t}|^{2}\right]+2\,\kappa^{2}\,|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|^{2}.$
Therefore we have
(3.4) $\frac{{\rm d}}{{\rm d}t}{\mathbb{E}}\left[|\overline{X}_{t}-\overline{Z}_{t}|^{2}\right]$ $\leq(-2\,\lambda+\lambda\,\kappa\,+2\,\sigma^{2})\,{\mathbb{E}}\left[|\overline{X}_{t}-\overline{Z}_{t}|^{2}\right]$
$\quad\quad\quad+(\lambda\,\kappa\,+2\,\sigma^{2}\,\kappa^{2})\,|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|^{2}.$
For sufficiently small values of $\kappa$ and $\sigma$, one guarantees $2\,\lambda+\lambda\,\kappa\,+2\,\sigma^{2}<0$. Let us define $\theta:=2\,\lambda-\lambda\,\kappa\,-2\,\sigma^{2}>0$, and $\gamma:=\lambda\,\kappa\,+2\,\sigma^{2}\,\kappa^{2}$. Therefore, Grönwall’s inequality yields
(3.5) $\mathbb{E}\left[\,|\overline{X}_{t}-\overline{Z}_{t}|^{2}\,\right]\leq e^{-\theta\,t}\mathbb{E}\left[\,|\overline{X}_{0}-\overline{Z}_{0}|^{2}\,\right]+\gamma\int_{0}^{t}|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{s})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|^{2}\,e^{-\theta(t-s)}\,\mbox{\rm d}s.$
Using $\mathbb{W}_{2}(\mu,\nu)^{2}=\inf\limits_{(Y\sim\mu,\,Y^{\prime}\sim\nu)}\mathbb{E}[|Y-Y^{\prime}|^{2}]$, we can take the infimum first in the left-hand side of (3.5), then in the right-hand side, and obtain
$\mathbb{W}_{2}(\rho^{\alpha}_{t},\rho^{\alpha}_{*})^{2}\leq e^{-\theta\,t}\mathbb{W}_{2}(\rho^{\alpha}_{0},\rho^{\alpha}_{*})^{2}+\gamma\int_{0}^{t}|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{s})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|^{2}\,e^{-\theta(t-s)}\,\mbox{\rm d}s.$
We are left with the last term which we estimate as follows. Since $\rho^{*}_{\alpha}\in\mathscr{P}_{2,R}(\mathbb{R}^{d})$, we can use (2.9) and obtain
$|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|\leq L_{\mathfrak{m}}\mathbb{W}_{2}(\rho^{\alpha}_{t},\rho^{\alpha}_{*}).$
Subsequently, we have
$\mathbb{W}_{2}(\rho^{\alpha}_{t},\rho^{\alpha}_{*})^{2}$ $\leq e^{-\theta\,t}\mathbb{W}_{2}(\rho^{\alpha}_{0},\rho^{\alpha}_{*})^{2}+\gamma\,L_{\mathfrak{m}}^{2}\,\int_{0}^{t}\mathbb{W}_{2}(\rho^{\alpha}_{s},\rho^{\alpha}_{*})^{2}\,e^{-\theta(t-s)}\,\mbox{\rm d}s$
$\Rightarrow\;e^{\theta\,t}\mathbb{W}_{2}(\rho^{\alpha}_{s},\rho^{\alpha}_{*})^{2}$ $\leq\mathbb{W}_{2}(\rho^{\alpha}_{0},\rho^{\alpha}_{*})^{2}+\gamma\,L_{\mathfrak{m}}^{2}\,\int_{0}^{t}\mathbb{W}_{2}(\rho^{\alpha}_{s},\rho^{\alpha}_{*})^{2}\,e^{\theta\,s}\,\mbox{\rm d}s$
Let us introduce $u(t):=e^{\theta\,t}\mathbb{W}_{2}(\rho^{\alpha}_{t},\rho^{\alpha}_{*})^{2}$, $a:=\gamma\,L_{\mathfrak{m}}^{2}$. Then we have
$u(t)\leq u(0)+\int_{0}^{t}a\,u(s)\,\mbox{\rm d}s$
We can now apply the integral version [^1] Suppose $u(t)\leq u_{0}+\int_{0}^{t}a(s)u(s)\,\mbox{\rm d}s+\int_{0}^{t}b(s)\,\mbox{\rm d}s$ with $a\geq 0$. Then denoting by $A(t)=\int_{0}^{t}a(s)\,\mbox{\rm d}s$, the following inequality holds: $u(t)\leq u_{0}e^{A(t)}+\int_{0}^{t}b(s)\,e^{A(t)-A(s)}\,\mbox{\rm d}s$. of Grönwall’s inequality and obtain
$u(t)$ $\leq e^{a\,t}u(0)$
$\Rightarrow\;\mathbb{W}_{2}(\rho^{\alpha}_{t},\rho^{\alpha}_{*})^{2}$ $\leq e^{(a-\theta)t}\,\mathbb{W}_{2}(\rho^{\alpha}_{0},\rho^{\alpha}_{*})^{2}\quad\forall\,t\geq 0$
where we recall $\theta=2\,\lambda-\lambda\,\kappa\,-2\,\sigma^{2}>0$ and $a=\gamma\,L_{\mathfrak{m}}^{2}=\kappa\,(\lambda+2\,\sigma^{2}\,\kappa)\,L_{\mathfrak{m}}^{2}$. For sufficiently small $\kappa$, we have $a-\theta<0$, and in (3.1) we have $\Theta:=\theta-a$.
To prove uniqueness of the invariant measure, it is sufficient to consider to dynamics governed by (3.3) but starting from two invariant measures $\rho^{\alpha}_{*,1},\rho^{\alpha}_{*,2}$. Subsequently, the contraction result in (3.1) yields
$\mathbb{W}_{2}(\rho^{\alpha}_{*,1},\rho^{\alpha}_{*,2})^{2}\leq e^{-\Theta\,t}\,\mathbb{W}_{2}(\rho^{\alpha}_{*,1},\rho^{\alpha}_{*,2})^{2}\quad\xrightarrow[t\to\infty]{}0$
which leads to a contradiction unless $\rho^{\alpha}_{*,1}=\rho^{\alpha}_{*,2}$, hence uniqueness. ∎
Remark 3.6.
We shall discuss the precise meaning of smallness of the parameters ensuring the validity of Theorem 3.3. As can be seen from its proof, the parameters need to verify the following two inequalities
(3.6) $\theta>a\;\Leftrightarrow 2\,\lambda-\lambda\,\kappa\,-2\,\sigma^{2}>\kappa\,(\lambda+2\,\sigma^{2}\,\kappa)\,L_{\mathfrak{m}}^{2}$
and $\theta>0\;\Leftrightarrow 2\,\lambda-\lambda\,\kappa\,-2\,\sigma^{2}>0$
The right-hand side of the first inequality is $\mathcal{O}(\kappa)$ which means that it can be made arbitrarily small, hence ensuring the validity of the inequality. In order to obtain an example of configuration ensuring the latter inequality, let us choose $\lambda>2\,\sigma^{2}$, and recall $\kappa\in(0,1)$. Therefore we have $\theta=2\,\lambda-\lambda\,\kappa\,-2\,\sigma^{2}>0$. So we need $\kappa\,(\lambda+2\,\sigma^{2}\,\kappa)\,L_{\mathfrak{m}}^{2}<2\,\lambda-\lambda\,\kappa\,-2\,\sigma^{2}$. A sufficient condition for Theorem 3.3 to hold is when
$\kappa\,(\lambda+2\,\sigma^{2}\,\kappa)\,L_{\mathfrak{m}}^{2}<2\lambda\,\kappa\,L_{\mathfrak{m}}^{2}<2\,\lambda-\lambda\,\kappa\,-2\,\sigma^{2},$
that is $2\lambda\,\kappa\,L_{\mathfrak{m}}^{2}+\lambda\,\kappa=\kappa\,\lambda(2L_{\mathfrak{m}}^{2}+1)<2\lambda-2\sigma^{2}$, thus it suffices to have $\kappa<\frac{2\lambda-2\sigma^{2}}{\lambda(2L_{\mathfrak{m}}^{2}+1)}$.
Conclusion: Taking into account Remark 2.7 and Remark 3.2, an example of sufficient conditions guaranteeing existence, uniqueness, and $\mathbb{W}_{2}$–exponential contraction is obtained when
$\delta=1,\quad\lambda>\max\{16\,\sigma^{2}\;,\;4\}\quad\text{ and }\quad 0<\kappa<\min\left\{\frac{2\lambda-2\sigma^{2}}{\lambda(2L_{\mathfrak{m}}^{2}+1)}\;,\;\frac{1}{2(1+b_{2})}\right\}$
where we recall $b_{2}$ is defined in (2.2), and $L_{\mathfrak{m}}$ is in Lemma 2.8.
3.3. Uniform second moment bound
The next result guarantees that the invariant measure has a uniform-in-$\alpha$ finite second moment. This will be important when studying the asymptotic $\alpha\to\infty$.
Theorem 3.7.
@@ 3.3. Uniform second moment bound
In the situation of Theorem 3.3, the unique invariant probability measure $\rho^{\alpha}_{*}$ to the rescaled CBO (1.6) satisfies
@@
Recall $\rho_{t}^{\alpha}:=\operatorname{Law}(\overline{X}_{t})$ the law of (1.6). According to Theorem 3.3, one has
$\mathbb{W}_{2}(\rho_{t}^{\alpha},\rho_{*}^{\alpha})^{2}\;\leq\;\mathbb{W}_{2}(\rho_{0}^{\alpha},\rho_{*}^{\alpha})^{2}\;e^{-\theta\,t}\quad\forall\,t\geq 0,\,\rho_{0}^{\alpha}\in\mathscr{P}_{2}(\mathbb{R}^{d})$
where $\theta$ is a positive constant. This implies in particular: for any $|\varphi(x)|\leq C(1+|x|^{2})$ it holds that
@@ 3.4. Harnack’s inequality
(3.10) $A=(a^{ij})_{i,j},\quad\text{ such that }\quad a^{ij}(x,\mu)=\frac{\sigma^{2}}{2}(\delta+|(x-\kappa\,\mathfrak{m}_{\alpha}(\mu))_{i}|)^{2}\delta_{ij}$
and $b=(b^{i})_{i},\quad\text{ such that }\quad b^{i}(x,\mu)=-\lambda(x-\kappa\,\mathfrak{m}_{\alpha}(\mu))_{i}$
$\delta_{ij}$ being the Kronecker symbol whose value is 0 whenever $i\neq j$ (not to be confused with the positive constant $\delta$). Moreover, the matrix $A$ is positive–definite because it is a diagonal matrix whose diagonal entries satisfy $\frac{\sigma^{2}}{2}(\delta+|(x-\kappa\,\mathfrak{m}_{\alpha}(\mu))_{i}|)^{2}\geq\frac{(\sigma\delta)^{2}}{2}>0$. In particular, when evaluated in the invariant probability measure $\rho_{*}^{\alpha}$, one obtains smooth coefficients (in $\mathcal{C}^{\infty}({\mathbb{R}}^{d})$)
@@
This identification is possible since the dependence of $A$ and $b$ on the measure is non-local – that is, it involves a quantity of the form $\int_{{\mathbb{R}}^{d}}K(y)\mbox{\rm d}\mu(y)$ which is moreover constant in $x$ (since the kernel $K$ does not involve xx)–, rather than a pointwise (local) evaluation of (the density of) the measure $\mu$. In other words, regularity of (the density of) the measure $\mu$ does not influence regularity of the coefficients $A$ and $b$.
@@
The following result is well-known and can be found for example in [5].
@@
First observe that the coefficients $A_{\nu}$ and $b_{\nu}$ are smooth, and $A_{\nu}$ is positive–definite thanks to $\delta>0$ in the diffusion. Then, existence and regularity of the density is [5, Corollary 1.6.7]. Harnack’s inequality is [5, Corollary 1.7.2]. ∎
@@
Existence and uniqueness of an invariant probability measure $\rho_{*}^{\alpha}$ is obtained in [32, Proposition 3.4 & Corollary 3.8], that is, $\rho_{*}^{\alpha}$ is the unique solution to the (non-linear) equation $L_{A,b,\mu}^{*}\mu=0$ in the class of probability measures. Moreover, (3.7) ensures that $\rho_{*}^{\alpha}$ has a finite second moment. Therefore, we can consider the diffusion (linear) operator $L_{A,b,*}$ and we are in the setting of Theorem 3.8 which gives the desired result when we choose the solution to $L_{A,b,*}^{*}\mu=0$ within the probability measures. Indeed in this case, the unique probability measure $\mu$ satisfying $L_{A,b,*}^{*}\mu=0$ is $\rho_{*}^{\alpha}$. Finally, the sets $U$ and $K$ in Theorem 3.8 are chosen as the closed balls $B_{2r}(x_{0})$ and $B_{r}(x_{0})$ respectively. ∎
@@
4. Laplace’s principle and global convergence
We are now ready to state and prove our main results concerning the limit $\alpha\to\infty$. We shall first treat the case where the function $f(\cdot)$ to be minimized has a unique minimizer in §4.1. Then, and although CBO methods are primarily designed for objective functions with a single minimizer, we are able to treat in §4.2 the more general situation where $f(\cdot)$ has multiple minimizing regions.
Let us recall that our results hold without any prior assumptions on the initial distribution $\rho_{0}\in\mathscr{P}_{4}(\mathbb{R}^{d})$ of (1.6). In particular, we do not assume that a global minimizer of $f(\cdot)$ is contained within the support of $\rho_{0}$. This absence of conditions on the initial distribution is precisely what ensures the global nature of the convergence.